### v2

# SONICS: Synthetic Or Not - Identifying Counterfeit Songs

Abstract
The recent surge in AI-generated songs presents exciting possibilities and challenges. While these tools democratize music creation, they also necessitate the ability to distinguish between human-composed and AI-generated songs for safeguarding artistic integrity and content curation. Existing research and datasets in fake song detection only focus on singing voice deepfake detection (SVDD), where the vocals are AI-generated but the instrumental music is sourced from real songs. However, this approach is inadequate for contemporary end-to-end AI-generated songs where all components (vocals, lyrics, music, and style) could be AI-generated. Additionally, existing datasets lack lyrics-music diversity, long-duration songs, and open fake songs. To address these gaps, we introduce SONICS, a novel dataset for end-to-end Synthetic Song Detection (SSD), comprising over 97k songs with over 49k synthetic songs from popular platforms like Suno and Udio. Furthermore, we highlight the importance of modeling long-range temporal dependencies in songs for effective authenticity detection, an aspect overlooked in existing methods. To capture these patterns, we propose a novel model, SpecTTTra, that is up to 3 times faster and 6 times more memory efficient compared to popular CNN and Transformer-based models while maintaining competitive performance. Finally, we offer both AI-based and Human evaluation benchmarks, addressing another deficiency in current research.
**footnotetext: Equal contribution22footnotetext: Code and Data will be available at https://github.com/awsaf49/sonics
1 Introduction
The rapid advancements in AI-generated music present a substantial threat to the music industry, potentially reducing the demand for professional musicians and stifling new talent development [19, 5]. To preserve the unique value of human creativity, it is crucial to develop robust methods for detecting AI-generated music, ensuring a fair and vibrant creative ecosystem.
Singing Voice Synthesis (SVS) [16] and Singing Voice Conversion (SVC) [11] have recently achieved significant progress, enabling the creation of synthetic singing voices that closely mimic real singers’ styles. When combined with instrumental music from real songs, these synthetic voices can produce convincing counterfeit songs. Although related to synthetic speech detection, detecting fake songs is particularly challenging due to the unique rhythmic patterns and artistic vocal traits of singing. To address this, researchers have turned their attention to Singing Voice Deepfake Detection (SVDD) [30, 34, 33]. However, current methods relying on datasets composed of SVS and SVC-generated songs face several limitations. These datasets are bound to use instrumental music from real songs, leading to artifacts like the "Karaoke effect" (volume discrepancies between music and vocals) and limited music-lyrics diversity. Moreover, existing methods often overlook the long-context temporal relationships inherent in songs, such as repeated verses, music, rhythm, and emotional dynamics, which are critical for effective detection. The short duration of songs in current datasets further hampers the use of these patterns. Additionally, copyright restrictions on some existing datasets limit the public availability of generated fake songs, hindering broader usage. Furthermore, the SVDD task requires separate tools for voice identification and separation during data processing [30, 34], increasing computational overhead.
Recently, platforms like Suno [24] and Udio [6] have gained significant traction on social media. They can synthesize not only vocals but also entire songs, including synthetic music, styles, and lyrics, further complicating the situation. Due to their end-to-end nature, these fake songs differ significantly from those generated by SVS and SVC methods, rendering existing SVDD methods and datasets inadequate for detecting them. This necessitates an urgent need for a detection system capable of identifying these end-to-end fake songs.
To address these shortcomings, we introduce SONICS, a large-scale dataset comprising over 97k songs, including more than 49k end-to-end synthetic songs generated by Suno and Udio, alongside 48k real songs collected from YouTube. With an average duration of 145 seconds, the SONICS dataset supports the use of long-context relationships in songs for accurate fake song detection. Our dataset also addresses the music-lyrics diversity issue by including fake songs generated with both real and synthetic lyrics and music. Additionally, the fake songs in our dataset are free from copyright issues, as they are generated using paid subscriptions. A unique feature of SONICS is that it includes the text lyrics of songs, which can aid future research.
Despite the availability of long songs in our dataset, utilizing long-context relationships presents additional challenges. For instance, CNN-based models struggle to capture long-range dependencies due to their inherently local receptive fields. While Transformer-based models can capture these dependencies with global attention, they are computationally expensive with longer audio inputs. To mitigate this trade-off, we introduce Spectro-Temporal Tokens Transformer (SpecTTTra), which uses a Spectro-Temporal Tokenizer to significantly reduce computational costs while enjoying global attention. Our contributions are summarized as follows:
  • We introduce SONICS, a large-scale dataset for synthetic song detection that addresses the limitations of existing datasets, including the absence of end-to-end fake songs, limited music-lyrics diversity, short-duration songs, and open fake songs.
  • We highlight the importance of modeling long-context temporal relationships in songs, an aspect entirely overlooked in existing approaches.
  • We propose the SpecTTTra model, which effectively captures long-context temporal relationships in songs at a low computational cost.
  • We offer a human benchmark for the fake song detection task, a feature missing in previous works.
  • We also establish a traditional AI benchmark, evaluating models across data partitions for deeper insights.
[…]
End-To-End
Fake Songs
[…]
Text Lyrics
[…]
Song Style
[…]
Music Diversity
[…]
Open Fake Songs
[…]
Open Real Songs
[…]
Open-Source
Models
[…]
Language
[…]
Average Length
[…]
# Algorithms
[…]
# Speakers
[…]
# Real Songs
[…]
# Fake Songs
[…]
# Total Songs
[…]
FSD Chinese 216.00 5 60 200 450 650
[…]
CtrSVDD Multi (no En.) 4.87 14 164 32,312 188,486 220,798
[…]
Table 1: Comparison of Proposed and Existing Fake Song Datasets
2 Related Works
Synthetic Speech Detection: The domain of synthetic speech detection, closely tied to synthetic song detection through their shared audio modality, has been extensively explored due to advancements in voice conversion [37] and synthesis techniques [28]. These advancements have spurred the development of audio spoofing attacks on speaker verification systems and deepfake audio targeting human listeners [13]. Countermeasures include Light CNN (LCNN) with Max-Feature-Map activations [15], Transformer encoders with ResNet architectures [36], RawNet2 with sinc layers and GRU blocks [25], and heterogeneous graph attention networks [12]. However, the unique complexities of songs—such as rhythm, melody, and emotional nuance—present challenges that traditional speech detection methods are not equipped to handle [30, 34]. Thus, following CtrSVDD [33], we opted not to conduct similar experiments in our study.
Synthetic Song Detection: Synthetic song detection, a relatively newer and more complex challenge, has gained attention recently. In 2023, SingFake [34] introduced a dataset of counterfeit songs using Singing Voice Conversion (SVC), along with the task of Singing Voice Deepfake Detection (SVDD) and associated model benchmarks. Subsequent work [30, 33] combined Singing Voice Synthesis (SVS) with SVC to create phoneme-based songs, leading to specialized detection datasets. Methods in this area include convolutional networks for feature extraction followed by graph neural network classification [33], wav2vec2-based extraction coupled with graph neural networks [26], and Linear-Frequency Cepstral Coefficients (LFCC) used with ResNet18 models [35]. However, SVC and SVS-based datasets always retain original background music, leading to a detectable "Karaoke effect" artifact. Recent end-to-end fake songs by Suno [24] and Udio [6], which can produce divergent fake songs where all musical components (e.g., background music, styles, and lyrics) can be synthetic, presenting a severe detection challenge.
Long Audio Classification: Songs exhibit long-range temporal patterns, such as repeated verses and rhythms, setting them apart from speech [1]. Despite their potential to enhance detection performance, these patterns have been largely overlooked in existing methods [30, 34, 33]. Meanwhile, long audio classification remains an uncharted area in audio research. Although automatic speech recognition handles long audio data [14], it struggles with end-to-end processing of extended audio due to its high computational cost, often using sliding window techniques [9, 21] to manage costs. This further complicates the task of leveraging long-context features for synthetic song detection.
3 Methodology
Split Label Source Count
Train Full Fake Suno-v3.5 774
Udio-130 656
Half Fake Suno-v3.5 4558
Mostly Fake Suno-v3.5 11863
Udio-130 16172
Real YouTube 32686
Total 66709
Test Full Fake Suno-v2 100
Suno-v3 206
Suno-v3.5 77
Udio-130 53
Udio-32 203
Half Fake Suno-v2 309
Suno-v3 634
Suno-v3.5 396
Mostly Fake Suno-v2 1350
Suno-v3 2797
Suno-v3.5 1117
Udio-130 1566
Udio-32 3970
Real YouTube 13237
Total 26015
Table 2: Dataset Summary
Figure 1: End-to-end pipeline of Full Fake song generation process. Here a random combination of topic, genre, and mood is utilized to create fake lyrics and fake song styles by prompt engineering LLM GPT-4o. The generated lyrics and styles are utilized to create synthetic songs through audio-generative models from Suno and Udio.
Figure 2: End-to-end pipeline of Mostly Fake song generation process. Firstly, lyrics features are extracted using an LLM from real song lyrics, which in turn is utilized to create fake lyrics through a second call to the LLM GPT-4o. Secondly, style information is extracted from real songs using multimodal Generative model Gemini 1.5 Flash. Finally, the generated lyrics and styles are utilized to create synthetic songs through audio-generative models from Suno and Udio.
Figure 3: Proposed Spectro-Temporal Tokens Transformer (SpecTTTra) model. First, the input mel-spectrogram undergoes separate temporal and spectral slicing to generate corresponding clips, which are tokenized into temporal and spectral tokens using separate tokenizers. Next, separate positional embeddings are added to these tokens, which are then passed through a Transformer encoder. The resulting globally contextualized features are then average pooled and finally passed to the classifier.
3.1 SONICS Dataset
The development of a modern synthetic song detection system necessitates a dataset that meets several stringent criteria, which are conspicuously absent in existing music datasets. These criteria include: 1) songs where all components—not just vocals—can be AI-generated; 2) song lengths sufficient to capture long-term contextual relationships; 3) a diverse spectrum of music-lyrics combinations; and 4) a quantity of data substantial enough to serve as a generative model benchmark. Addressing these needs, we introduce the SONICS dataset, a comprehensive collection of end-to-end AI-generated songs produced using the latest audio generative models, spanning lengths from 32 to 240 seconds. The dataset encompasses an extensive array of music-lyrics styles, amassing a total of 97,164 songs. A detailed comparison of SONICS with existing datasets is presented in Table 1, clearly illustrating that datasets such as FSD [30], SingFake [34], and CtrSVDD [33] fall short of fulfilling all the outlined criteria when juxtaposed with SONICS. Additionally, a comprehensive distribution summary of the SONICS dataset is provided in Table 2.
3.1.1 Real Songs
The dataset’s Real songs segment comprises original compositions by human artists. This portion of the dataset was initially compiled by random sampling from the Genius Lyrics Dataset [10], which provides metadata including lyrics, titles, and artist information. Subsequently, leveraging this metadata, an API-based dynamic search was conducted on YouTube to retrieve the corresponding audio files [2]. This process yielded 48,090 original songs performed by 9,096 artists.
3.1.2 Fake Songs
The Fake songs subset consists of compositions generated by AI models guided by two primary text inputs: 1) lyrics and 2) song-style (e.g., instruments, genre, vocal type). Based on the nature of these inputs, the Fake songs are categorized into three distinct groups: i) Full Fake (FF)—featuring both AI-generated lyrics and song-style; ii) Mostly Fake (MF)—where lyrics are AI-generated based on real lyrics features and song-style is derived from real songs; and iii) Half Fake (HF)—where the lyrics are directly sourced from real songs, with song-style also extracted from real compositions.
To generate FF songs, we curated a rich set of metadata, including 57 broad topics (e.g., friendship, betrayal), 292 specific topics (e.g., star trek, pokemon), 49 music genres (e.g., rock, metal), and 72 moods (e.g., calm, angry). Random combinations of these elements were used to generate final lyrics and styles via a Large Language Model (LLM). The complete pipeline for FF song generation is illustrated in Fig. 1. In contrast, the creation of MF and HF songs involved extracting song- styles from real songs using the multimodal LLM Gemini 1.5 Flash, which uniquely supports audio input. For MF songs, lyrics were generated using the GPT-4o LLM, which extracts features from existing real song lyrics to produce new lyrics with a similar distribution, avoiding direct replication of existing content. The pipeline for generating MF songs is detailed in Fig. 2. The HF songs, differing only in the use of real lyrics directly sourced from existing songs, follow a similar process. It is noteworthy that, Udio models cannot generate songs for lyrics from real songs; thus, HF songs contain only Suno models. Using this curated metadata, a variety of audio generative models were employed to produce a total of 49,074 Fake songs, comprising 6,132 HF, 40,769 MF, and 2,173 FF songs. The prompt templates used in LLM to extract lyrics feature and create lyrics along with song- styles can be found in the appendix.
For generating these end-to-end songs, we utilized the only two currently available audio generative model families: Suno [24] and Udio [6]. Specifically, we incorporated three variants from Suno—Suno v2, v3, and v3.5. Suno v2, the earliest iteration, generates songs up to 80 seconds (sec.) in length. Suno v3 extends this capability to 120 sec, offering enhanced audio quality and broader stylistic support. The latest iteration, Suno v3.5, produces songs up to 240 sec, featuring improved song structure and vocal flow. From the Udio family, we used Udio-32 and Udio-130, capable of generating songs of 32 and 130 sec, respectively. It is noteworthy that, both the Suno and Udio models generate two songs per (lyrics, style) input.
3.2 Model
The architecture of an audio classification model is influenced by how the input audio is processed for feature extraction. In our benchmark, we primarily use Spectrograms more specifically Mel-spectrogram due to their versatile usage and effective performance across various audio processing tasks [8, 21, 22, 20, 23]. Given the 2D image-like nature of Spectrogram, image classifier models are used for classifying the audio. Given that image classification models can broadly be divided into two categories: CNN and Transformer-based architectures, in our work, we employ popular models from both categories, such as ConvNeXt [18], ViT [7], for benchmarking. However, these models encounter challenges when dealing with the inherent long-range temporal dependencies in songs. Specifically, ConvNeXt struggles to capture long-range dependencies due to its local receptive field, a characteristic inherited from CNNs. On the other hand, while ViT is capable of capturing long-range dependencies, it incurs significant computational costs as the number of patches/tokens rapidly increases with longer audio inputs. To address this issue, we introduce SpecTTTra, which utilizes global attention similar to ViT but employs a Spectro-Temporal Tokenizer to reduce the number of tokens considerably, thereby lowering computational costs and enhancing efficiency significantly.
3.2.1 Spectro-Temporal Tokens Transformer
As illustrated in Fig. 3, the proposed SpecTTTra model begins by applying temporal and spectral slicing to copies of the input spectrogram, generating distinct temporal and spectral clips (or patches). These clips are then processed by the Spectro-Temporal Tokenizer, where temporal clips are embedded as temporal tokens $\langle t_{i}\rangle$ and spectral clips as spectral tokens $\langle f_{i}\rangle$, using separate tokenizers. Each token is represented as a vector of shape $\left(1,embed\_dim\right)$. Subsequently, separate positional embeddings are added to the temporal and spectral tokens, as there is no positional relationship between these two types.
The positionally aware tokens are then fed into a ViT-like Transformer encoder, where they are contextualized with one another through global attention. This process enables the temporal tokens to become aware of both other temporal tokens and spectral tokens, and vice versa for spectral tokens. These globally aware features, now of shape $\left(n\_tokens,embed\_dim\right)$, are average-pooled across the $n\_tokens$ dimension to aggregate the temporal and spectral information. Finally, the accumulated features are passed to a classifier, where they are classified as either real or fake songs.
It is important to note that while previous work has attempted to utilize both spectral and temporal information, these approaches come with significant limitations. For instance, [31] focuses solely on temporal tokens, neglecting the rich spectral information. On the other hand, methods like [32, 23] employ separate attentions for spectral and temporal information but with ViT-like tokens, which, as discussed in the following section, become highly inefficient and computationally expensive as they dramatically increase with longer audio inputs. In contrast, our approach disentangles spectral and temporal information at the tokenization level and then contextualizes them with attention. This results in a more efficient method compared to existing approaches.
3.2.2 Spectro-Temporal Tokenization
Traditional Vision Transformers (ViTs) generate grids of patches by simultaneously dividing 2D spectrogram across both temporal and spectral dimensions. If $T$ and $F$ denote the temporal and spectral dimensions of the spectrogram, respectively, and $p$ is the patch size, then the number of patches (or tokens) generated by a ViT ($N_{\nu}$) can be expressed as:
$N_{\nu}=\left(\frac{F}{p}\right)\times\left(\frac{T}{p}\right)=\left(\frac{FT}{p^{2}}\right)$ (1)
This equation highlights the rapid growth in the number of patches as the temporal dimension $T$ increases, leading to a corresponding exponential rise in computational cost. For example, with a spectrogram of size $F=128$ (mel bands) and short audio with $T=128$ (5 sec.), ViT generates $N_{\nu}=64$ tokens (for $p=16$). However, for longer audio with $T=3744$ (120 sec.), the number of tokens surges to $N_{\nu}=1872$, nearly 30 times more than for the shorter audio. Since the computational cost for global attention scales quadratically with the number of tokens, this makes ViTs impractical for long audio classification tasks.
In contrast, the proposed SpecTTTra model adopts an alternative strategy by independently slicing across temporal and spectral dimensions to create separate temporal and spectral patches, as illustrated in Fig. 3. Unlike ViTs, where each patch only has access to limited spectral and temporal information due to simultaneous slicing across both dimensions, SpecTTTra ensures that each temporal patch has access to all spectral components and vice versa. This design leverages the observation that meaningful correlations can exist between distant temporal clips (e.g., between $|t_{0}|$ and $|t_{4}|$, capturing repeated song verses) or between distinct spectral clips (e.g., between $|f_{0}|$ and $|f_{2}|$, differentiating real from fake songs). If $t$ and $f$ represent the sizes of the temporal and spectral clips, respectively, then the number of patches (or tokens) generated by SpecTTTra, denoted as $N_{\psi}$, is given by:
$N_{\psi}$ $=N_{spectral}+N_{temporal}=\left(\frac{F}{f}\right)+\left(\frac{T}{t}\right)$ (2)
Due to the additive nature of this equation, the number of tokens in SpecTTTra grows much more slowly compared to ViTs, thereby keeping computational costs manageable. For instance, using the same parameters as before and setting $t=7$ and $f=5$ in SpecTTTra, the number of tokens for short audio with $T=128$ is $N_{\psi}=43$, and for long audio with $T=3744$, it increases to $N_{\psi}=560$, which is approximately 3.4 times fewer than in the ViT model. This substantial reduction in token count makes SpecTTTra significantly more computationally efficient. We further explore three variants of SpecTTTra, differentiated by the sizes of their temporal ($t$) and spectral ($f$) patches: SpecTTTra-$\alpha$ ($f=1$, $t=3$), SpecTTTra-$\beta$ ($f=3$, $t=5$), and SpecTTTra-$\gamma$ ($f=5$, $t=7$). Fig. 4 illustrates the rate at which the number of tokens increases for different SpecTTTra variants and ViTs as the audio length (number of time frames) grows.
Figure 4: Comparison of the number of tokens generated by the ViT model and the three SpecTTTra variants ($\alpha$, $\beta$, and $\gamma$) as a function of the number of time frames.
The resulting spectral and temporal clips are processed by the Spectro-Temporal Tokenizer (STT) to create spectral and temporal tokens, respectively. The proposed STT block consists of separate Spectral and Temporal Tokenizers, which share identical architecture but differ in their objectives based on their input. These tokenizers use a Linear layer to map the spectral/temporal clips to spectral/temporal tokens, followed by GELU activation, addition of learnable positional embeddings similar to ViT, and finally, layer normalization. For efficiency, the slicing and tokenization operations are merged using 1D convolution. Given an input spectrogram $x\in\mathbb{R}^{F\times T}$, where $T$ and $F$ denote the temporal and spectral dimensions of the input spectrogram, the mathematical expression for tokenization is as follows:
[…]
$\langle t\rangle$ $=\text{LayerNorm}\left(\text{GELU}\left(x^{\top}_{t}\right)+\hat{p}_{t}\right)$
[…]
$\langle f\rangle$ $=\text{LayerNorm}\left(\text{GELU}\left(x^{\top}_{f}\right)+\hat{p}_{f}\right)$
Here, $\langle t\rangle\in\mathbb{R}^{\frac{T}{t}\times D}$ and $\langle f\rangle\in\mathbb{R}^{\frac{F}{f}\times D}$ denote the temporal and spectral tokens, where $t$ and $f$ are the temporal and spectral clip sizes, respectively, and $D$ is the token embedding dimension. The parameters $c_{i}$, $c_{o}$, $k$, and $s$ represent the input channels, output channels, kernel size, and stride size of the CNN layers. Additionally, $\hat{p}$ and $\top$ denote the learnable positional embeddings and Transpose operation. Further implementation details can be found in the Appendix.
Len. (sec) Model
Algorithm
Singer
Fake Type
Overall
Suno
v3.5
Suno
v3
Suno
v2
Udio
32
Udio
130
Seen Unseen
Half
Fake
Mostly
Fake
Full
Fake
F1 Sens. Spec.
5 ConvNeXt 0.99 0.99 0.62 0.62 0.99 0.99 0.99 0.90 0.82 0.80 0.90 0.82 0.98
ViT 0.98 0.95 0.79 0.57 0.86 0.79 0.79 0.92 0.78 0.76 0.79 0.80 0.79
SpecTTTra-$\gamma$ 0.99 0.98 0.51 0.10 0.99 0.98 0.97 0.87 0.61 0.62 0.76 0.63 0.98
SpecTTTra-$\beta$ 0.99 0.98 0.61 0.18 0.99 0.95 0.94 0.89 0.66 0.66 0.78 0.69 0.94
SpecTTTra-$\alpha$ 0.99 0.99 0.68 0.26 0.99 0.93 0.92 0.91 0.69 0.70 0.80 0.71 0.92
120 ConvNeXt 0.99 0.99 0.77 0.95 1.0 0.98 0.98 0.94 0.95 0.93
0.96
0.95 0.98
ViT 1.00 0.99 0.82 0.53 0.99 0.99 0.98 0.95 0.80 0.80 0.89 0.82 0.98
SpecTTTra-$\gamma$ 0.99 0.99 0.83 0.59 0.99 0.99 0.98 0.95 0.82 0.81 0.91 0.83 0.98
SpecTTTra-$\beta$ 1.00 0.99 0.87 0.62 0.99 0.99 0.99 0.96 0.84 0.82 0.92 0.86 0.99
SpecTTTra-$\alpha$ 1.00 0.99 0.88 0.71 0.99 0.99 0.99 0.98 0.89 0.87 0.94 0.91 0.99
Table 3: AI Benchmark
4 Experiments and Results
4.1 Dataset
We conduct all experiments using the proposed SONICS dataset, which is divided into train, valid, and test sets. To ensure comprehensive evaluation, the valid and test sets include cases with unseen algorithms (e.g., Suno v2, Suno v3, Udio 32) and new singers. We also prevent data leakage by ensuring that song pairs from the same (lyrics, style) inputs are exclusively in either the training or valid-test sets, not in both. The distribution of the train and test sets is shown in Table 2, with the valid set distribution available in the Appendix.
4.2 Implementation Details
We conduct model training on an NVIDIA A6000 GPU with 48GB RAM, using WandB [3] for tracking. Generating the SONICS dataset costs \$1,055, allocated across GPT-4o (\$405), Gemini 1.5 Flash (\$80), Suno (\$390), and Udio (\$180). We generate spectrograms with a 16,000 Hz sample rate, n_fft = win_length = 2048, hop_length = 512, and n_mels = 128, yielding a $128\times 128$ spectrogram for 5 sec and $128\times 3744$ for 120 sec audio. Any song shorter than input length is zero-padded randomly, while for longer songs, a random crop is used. However, during the test, to maintain determinism, padding is done on the right side and cropped segments are taken from the middle. We use small and tiny variants of ViT (patch size = 16) and ConvNeXt from the timm [29] library. In SpecTTTra, we use the same model configuration as ViT-small. It is noteworthy that we trained all models from scratch for fair comparison. We train for 50 epochs using Binary Cross-Entropy loss with 0.02 label smoothing and a batch size of 128. Optimization is performed with AdamW and a cosine learning rate scheduler from timm, including a 5-epoch warm-up and an initial learning rate of 0.0005. While existing methods [34, 33, 30] typically use Equal Error Rate (EER) as a metric, we prioritize the F1 score (binary average, threshold = 0.5) due to EER’s susceptibility to class imbalance. We also evaluate Sensitivity (Sens.) and Specificity (Spec.) to assess performance across fake and real classes. Further details can be found in the Appendix.
4.3 Benchmarks
4.3.1 AI Benchmark
The comparative analysis of the proposed SpecTTTra models against ViT and ConvNeXt models is presented in Table 3. The results reveal a significant performance gain (6% for ConvNeXt, 10% for ViT, and 14% for SpecTTTra-$\alpha$) in the Overall F1 score when using long songs. This finding substantiates our claim that leveraging long-context information is crucial for enhancing fake song detection. Furthermore, the advantage of longer audio duration is more prevalent in transformer-based models (ViT or SpecTTTra variants) compared to CNN models (ConvNeXt). This can be attributed to the global attention mechanism in transformer models, which effectively captures long-range dependencies within the input data. However, despite the absence of global attention, ConvNeXt outperforms both ViT and SpecTTTra models in both short and long audio scenarios. We hypothesize that this is due to the inherent inductive biases present in CNNs, which are lacking in transformers, leading the latter to require larger datasets to achieve comparable performance [17, 7]. Notably, the proposed SpecTTTra-$\alpha$, while trailing ConvNeXt by 10% in the F1 score for short audio, narrows this gap to 4% for long audio, highlighting its potential in long-context detection tasks. Another intriguing observation is the performance of ViT, which, despite its large number of tokens (or patches), is outperformed by all SpecTTTra variants in terms of overall F1 score for long audio, solidifying SpecTTTra’s effectiveness. This could be due to an overload of redundant information from ViT’s numerous patches, which may not contribute effectively to the detection task. Moreover, an interesting pattern can be observed across all models – real songs are more easily identified than fake ones, as indicated by higher specificity and lower sensitivity scores.
Diving deeper into different partitions of test data, we observe that all detection models achieve better performance on seen algorithms (Suno v3.5 and Udio 130) compared to unseen ones (Suno v2, Suno v3, and Udio 32). Particularly, they struggle more with the Udio algorithms, with the most pronounced difficulty observed for Udio 32. However, ConvNeXt performs relatively well in detecting the Udio 32 algorithm, achieving a sensitivity of 95%, possibly due to the CNN’s strong locality. Interestingly, despite being an unseen algorithm, the detectors perform comparably well on Suno v3 as they do on the seen Suno v3.5 algorithm, suggesting a possible algorithmic similarity between the two. On the other hand, for short audio samples, the models perform better on songs with seen speakers than those with unseen speakers, a gap that diminishes when longer audio is used. Finally, in fake type partitions, all detectors excel in detecting HF songs, which can be linked to the exclusive presence of Suno algorithms, where the detectors generally perform better compared to Udio algorithms. Among MF and FF songs, the models show slightly lower performance on FF songs, which might be due to the diverse unseen combination of topics, moods, and genres that differ more significantly than other fake songs.
4.3.2 Human-AI Benchmark
To evaluate Human performance in comparison to AI-based models, we selected a subset of 520 samples from our large test data. This evaluation employed a dynamic scoring system, similar to LMSYS [4], allowing public participation and live leaderboard updates, which will be made publicly available upon acceptance of this paper. Three human participants were involved in this benchmark, with their performance summarized in Table 4. In contrast to the AI benchmark using short (5 sec) or long (120 sec) audio samples, this human benchmark employed 25 sec clips. This choice stems from the observation that short clips hinder human identification due to subtle inaudible artifacts easily detected by AI, while longer clips don’t necessarily improve human performance due to the difficulty in leveraging long-range temporal dependencies.
As shown in Table 4, AI-based methods consistently outperform human participants across all test partitions. However, both humans and AI models struggled most with Udio algorithms, particularly Udio-32, where human sensitivity dropped to 23.4%. Conversely, Suno algorithms, especially Suno-v3.5, were easier to detect, with a human sensitivity of 81.7%. This mirrors the findings in the AI benchmark, where models demonstrated higher specificity than sensitivity, indicating greater ease in identifying real songs compared to fake ones. Further analysis revealed distinct characteristics within real and fake song audio patterns. For instance, Suno algorithms often produced synthetic or mechanical-sounding vocals, while Udio-32 algorithm occasionally created the "Karaoke effect." Furthermore, Udio algorithms demonstrated the ability to create songs with multiple voices and higher notes, a feature absent in Suno algorithms. On the other hand, real songs exhibited unique elements like a wide range of notes, diverse timbre, complex rhythms, and clear vocals with accented pronunciation. They also featured distinctive sounds like flutes and finger snaps, which are challenging for fake songs to replicate.
Partition
ConvNeXt
ViT
SpecTTTra-$\alpha$
SpecTTTra-$\beta$
SpecTTTra-$\gamma$
Human
Algorithm (Sens.) Suno v2 0.65 0.80 0.72 0.64 0.54 0.69
Suno v3 0.99 0.96 0.99 0.98 0.98 0.75
Suno v3.5 0.99 0.98 0.99 0.99 0.99 0.82
Udio 32 0.67 0.56 0.33 0.23 0.18 0.23
Udio 130 0.99 0.87 0.99 0.99 0.99 0.55
Fake Type (Sens.) Half Fake 0.91 0.93 0.92 0.90 0.88 0.71
Mostly Fake 0.84 0.79 0.72 0.69 0.64 0.66
Full Fake 0.83 0.77 0.72 0.68 0.64 0.63
Real (Spec.) 0.98 0.80 0.94 0.95 0.97 0.78
Fake (Sens.) 0.85 0.82 0.75 0.72 0.66 0.66
Overall (F1) 0.92 0.82 0.83 0.80 0.78 0.71
Table 4: Human-AI Benchmark
4.3.3 Efficiency Benchmark
To comprehensively evaluate the efficiency of the proposed SpecTTTra model alongside other methods, we measure various metrics across different song lengths using a P100 16GB GPU. The metrics considered include Speed (A/S $\rightarrow$ Audio per Second), Floating Point Operations (FLOPs), GPU Memory Consumption (Mem.) during the forward pass with a batch size of 14, activation count (# Act.), and parameter count (# Param.). The results are summarized in Table 5. Our analysis reveals that while ViT is the fastest model for 5-sec songs, it becomes the slowest for 120-sec songs and exhibits significant memory consumption, rendering it impractical for longer sequences. However, ViT remains the most efficient in terms of parameter count across both short and long songs. On the other hand, ConvNeXt, despite its strong detection performance, is the most resource-intensive model in terms of memory usage, FLOPs, and parameter count. In contrast, the proposed SpecTTTra model variants demonstrate remarkable efficiency, particularly in longer songs. Although they may not excel in short audio segments, the SpecTTTra-$\gamma$ variant is nearly 3 times faster than ViT and 2 times faster than ConvNeXt for 120-sec songs. Moreover, it consumes memory 3 times less than ViT and 6 times less than ConvNeXt, highlighting its superior efficiency. Thus, while SpecTTTra secures the 2nd position in the detection benchmark, it offers the best trade-off between detection and efficiency performance, making it the most balanced model overall.
Len.
Model
Speed
(A/S)
FLOPs
Mem.
# Act.
# Param.
5 ConvNeXt 141 1.5 0.5 4 28
ViT 168 1.1 0.2 2 17
SpecTTTra-$\gamma$ 164 0.7 0.1 1 17
SpecTTTra-$\beta$ 162 1.1 0.2 2 17
SpecTTTra-$\alpha$ 153 2.8 0.6 6 17
120 ConvNeXt 39 43.1 13.6 129 28
ViT 34 31.7 6.2 67 17
SpecTTTra-$\gamma$ 97 10.1 1.9 20 24
SpecTTTra-$\beta$ 80 13.9 2.6 29 21
SpecTTTra-$\alpha$ 48 23.6 4.6 50 19
Table 5: Efficiency Benchmark
5 Limitations and Future Work
The real songs in the proposed SONICS dataset are queried from YouTube dynamically using their title and artist name, which introduces the possibility of not retrieving the correct audio. Through manual analysis of random samples, we estimate that this issue affects approximately 0.5% of the dataset. Also, due to the copyright of the real songs, we are unable to make them public. However, our dataset contains YouTube links for each song, which can be used to download the song easily using API [2]. Another limitation arises from the fake songs generated by the Udio platform, which cannot include lyrics from real songs. This restricts comprehensive evaluation in the Half Fake songs category, as it only contains fake songs from the Suno platform. Furthermore, the dataset currently includes only English-language songs, which may restrict the scope of our evaluation. In future work, we plan to extend our dataset to multiple languages to enable a more comprehensive analysis. Additionally, our current benchmarks are based solely on Mel Spectrogram input. Moving forward, we aim to incorporate raw audio and explore other feature extraction methods, such as LFCC and MFCC, to enhance the robustness of our evaluations.
6 Conclusion
In this paper, we introduced SONICS, a comprehensive dataset for end-to-end synthetic song detection, addressing limitations in existing datasets, such as lack of music diversity, short duration, and most importantly, the absence of end-to-end AI-generated songs. Moreover, we proposed the SpecTTTra model, which efficiently captures long-range temporal relationships in songs, achieving competitive performance to existing popular models while reducing computational costs significantly. Through extensive experiments, we established both AI-based and human benchmarks, demonstrating the dataset’s effectiveness in advancing synthetic song detection research. Our work lays the foundation for improved methods for ensuring fair attribution and protecting the integrity of musical artistry.
7 Acknowledgment
We express our gratitude to the Suno and Udio platforms, which we used for generating end-to-end synthetic songs. We also appreciate the GPU credits provided by Jarvislabs.ai, which were partially used for training our model.
Appendix
Appendix A Dataset Statistics
Long Form Correlations: An example of the long context correlation can be observed on Fig. 5, where a real song would stay consistent throughout the repetition of the phrases, refrains, rythms, etc. In comparison, synthetic generation methods can have difficulty generating consistent refrains, due to the long form context that might be required to model this information. Therefore, a synthetic song detection model that can capture the nuances of long context correlations present in songs would be able to better differentiate real and synthetic songs utilizing this property.
Figure 5: Long Context Correlation
Validation Data: In Table 6, the distribution of the labels in validation data can be seen. For proper evaluation, there is are 104 Full Fake, 235 Half Fake and 1,934 Mostly Fake songs, cumulating into 2,273 Fake songs. In comparison, there are 2,167 Real songs taken from YouTube in the validation data.
Table 6: Valid Data Summary
Label Source Count
Full Fake Suno-v2 16
Suno-v3 37
Suno-v3.5 9
Udio-130 12
Udio-32 30
Half Fake Suno-v2 51
Suno-v3 117
Suno-v3.5 67
Mostly Fake Suno-v2 258
Suno-v3 494
Suno-v3.5 196
Udio-130 286
Udio-32 700
Real YouTube 2167
Total 4440
Song Duration: The distribution of the song duration across training and test sets is shown on Fig. 6. From the figure, it can be observed that aside from a few minor outliers, the distribution of fake and real song durations across both training and test sets are similar.
Figure 6: Duration distribution between Training and Testing sets
Embedding Space: The t-SNE plots presented in Fig. 7 illustrate the data distributions within the embedding space, generated using an EfficientNetB0 [27] encoder with Mel Spectrogram audio inputs.
  • Subfigure (a) shows a significant overlap between the training (Green) and testing (Orange) datasets. However, distinct clusters along the edges, predominantly composed of training data, suggest that the training set is more diverse than the test set. This diversity is observed despite the test set containing previously unseen algorithms and speaker scenarios.
  • Subfigure (b) indicates a substantial overlap between real (Red) and fake (Blue) songs, with a few regions exclusively occupied by fake data or outliers. This highlights the inherent challenge of synthetic song detection within this dataset.
  • Subfigure (c) shows that the embedding space is primarily occupied by Suno v3.5 (Green) and Udio 133 (Orange), which are seen algorithms. Despite Suno v2 (Blue), Suno v3 (Red), and Udio 32 (Violet) being unseen algorithms, their embeddings fall within regions covered by seen algorithms, demonstrating underlying similarities. A noteworthy observation is the considerable overlap between Suno v3 (Red) and Suno v3.5 (Green), which likely contributes to the similar detection performance in these algorithms, as indicated by Table 3.
  • Subfigure (d) depicts the distribution of fake types. Although Half Fake (Orange) songs form discernible clusters, Fully Fake (Red) songs exhibit a more scattered distribution, indicating a lack of common features. Conversely, Mostly Fake (Green) songs are spread broadly across the embedding space, reflecting their diverse characteristics.
Figure 7: Data distributions in the embedding space. This plot illustrates the clustering of different data categories such as target, algorithms, fake types, and data split.
Song Style: In order to imitate the song style observed in real songs, we had LLMs perform a stylistic analysis of the real song lyrics (prompt shown in Table 9). The wide variety of styles extracted from real songs are utilized to generate the synthetic songs. It is important to have the distribution of synthetic song style as close to the real song styles in the dataset, since it would be a better indicator of real life cases. A wide diversity of synthetic song style in training data is also required so that the models trained on this dataset do not rely on the distinct characteristics of any particular song style as a feature to detect synthetic songs. Additionally, it is also important to evaluate performance of the trained model on a testing set across a wide variety of song style as well, so that the generalization capability of the detection model across multiple styles is evaluated. A word cloud representation of the song styles that have been extracted from real songs and utilized to generate songs is show on Fig. 8. The figure indicates that there are diverse song styles across both training and testing sets.
Figure 8: Data distributions of song-style between train and test split in the word cloud
Figure 9: Song Arena on Huggingface
Figure 10: Song Arena Leaderboard on Huggingface
Appendix B Implementation Details
To provide a clear understanding of the proposed SpecTTTra model, we present below the PyTorch-like code for its core component, the Spectro-Temporal Tokenizer ($st\_tokenizer$). It is important to note that although the provided code is in a functional format for clarity, an object-oriented approach was used in the actual implementation.
{minted}
[bgcolor=bgcolor, fontsize=]python import torch import torch.nn as nn
def st_tokenizer(x, t_clip, f_clip, embed_dim): B, F, T = x.size() # Temporal tokens t_tokens = tokenizer( x, F, embed_dim, t_clip, T // t_clip, ) # Spectral tokens f_tokens = tokenizer( x.transpose(1, 2), T, embed_dim, f_clip, F // f_clip, ) # Spectro-Temporal tokens st_tokens = torch.cat( (t_tokens, f_tokens), dim=1 ) return st_tokens
def tokenizer( x, input_dim, token_dim, clip_size, n_clips ): # Slicing and Tokenization conv1d = nn.Conv1d( in_channels=input_dim, out_channels=token_dim, kernel_size=clip_size, stride=clip_size, bias=False, ) x = conv1d(x).gelu().transpose(1, 2) # Positional Embedding pos_embeds = nn.Parameter( torch.randn(1, n_clips, token_dim) * 0.02 ) x = x + pos_embeds # Layer Normalization x = nn.LayerNorm(token_dim, eps=1e-6)(x) return x
In this code, the $tokenizer$ function represents the spectral or temporal tokenizer used to embed spectral or temporal clips (patches) into tokens. Here, $f\_clip$ and $t\_clip$ denote the sizes of the spectral and temporal clips, respectively, while $embed\_dim$ signifies the feature dimension of each token. The dimensions $T$ and $F$ correspond to the temporal and spectral dimensions of the input spectrogram.
Appendix C Human-AI Benchmark
To evaluate the human performance of synthetic song detection, a Huggingface space is titled "Song Arena" (shown in Fig. 9) is created where users can evaluate whether a random song is synthetic or not from a small subset (520 samples in total) of the proposed dataset. A leaderboard (shown in Fig. 10) consisting of the detection rate of the samples belonging to each algorithm as well as generation methods (mostly fake, half fake and full fake songs - as outlined in Section 3.1.2) are also part of the space. Sensitivity (True Positive Rate) is used to evaluate the performance of synthetic song generation algorithms.
Appendix D Prompt Engineering
Half Fake: These songs are generated using lyrics and song style extracted from real songs. To extract the song style, the prompt template mentioned in Table 9 is used. This template extracts song style information such as vocal type, musical instruments, mood, etc.
Mostly Fake: These songs are generated similarly to Half Fake songs, except that the lyrics are AI-generated. The lyrics are created using an LLM (Large Language Model) with the prompt template shown in Table 8, where lyrics features are used as input. These lyrics features are also extracted from real song lyrics using the prompt template shown in Table 7. The use of lyrics features instead of direct lyrics prevents the LLM from copying the original content, encouraging it to generate lyrics with a similar distribution rather than duplicating them.
Full Fake: These songs are generated using AI-generated lyrics and song style, created using the prompt template provided in Table 10. In this process, the genre, topic, and mood were selected randomly from the lists provided below:
  • List of Genres: alternative, baroque, blues, bollywood, c-pop, celtic, christian rock, classical, country, crunk, dance, dancehall, disco, doom metal, electronic, folk, funk, fusion, gospel, gothic, grime, grunge, hard rock, heavy metal, hip hop, indie rock, j-pop, jazz, k-pop, lo-fi, lounge, metal, metalcore, new age, opera, orchestral, pop, pop rock, progressive metal, progressive rock, punk, r&b, rap, reggae, salsa, smooth jazz, soul, sufi, world music.
  • List of Moods: adventurous, ambivalent, amused, angry, anxious, apathetic, bittersweet, blissful, calm, carefree, cautious, chaotic, confident, confused, curious, desperate, determined, disenchanted, distracted, drained, dreamy, empathetic, enchanted, energetic, exhilarated, focused, forgiving, frustrated, gloomy, grateful, hateful, humble, inspired, introspective, jealous, joyful, liberated, lonely, loving, melancholic, mischievous, motivated, mournful, mysterious, nostalgic, optimistic, passionate, pensive, pessimistic, playful, powerless, proud, rebellious, regretful, reluctant, restless, romantic, sarcastic, satisfied, shocked, skeptical, submissive, sympathetic, tense, timid, trapped, uninspired, vengeful, vulnerable, whimsical, yearning, zealous.
  • List of Broad Topics: alien invasion, ancient civilizations, augmented reality, betrayal, childhood memories, climate change, cyber crime, dimensional portals, dreams and aspirations, dystopian future, empowerment, endangered species, extraterrestrial contact, family, fashion, financial struggles, first kiss, forgiveness, friendship, futuristic cities, generation gap, grief and loss, heartbreak, social media anxiety, interstellar travel, loneliness in a crowd, long-distance relationships, love, love at first sight, lunar colonization, nanotechnology, nature’s beauty, nostalgia, ocean exploration, overcoming adversity, pandemic experiences, parallel universes, political revolution, politics, quantum physics, reincarnation, never give up, road trip adventures, robotic emotions, save the planet, sibling rivalry, social influencers, social justice, space exploration, space tourism, survival in the wild, technology addiction, time capsules, time paradoxes, time travel, unconditional love, work-life balance.
  • List of Specific Topics: elon musk vs yann lecun (AI), yann lecun vs geoffrey hinton (AI), convolution vs transformer (AI), gan vs diffusion models (AI), tensorflow vs pytorch (AI), pytorch vs jax (AI), twilight zone (TV), star trek (TV), game of thrones (TV), breaking bad (TV), stranger things (TV), big bang theory (TV), friends (TV), simpsons (TV), house of cards (TV), how I met your mother (TV), the office - US (TV), sherlock (TV), avatar - the last airbender (TV), pokemon (anime), dragon ball z (anime), naruto (anime), one piece (anime), attack on titan (anime), my hero academia (anime), death note (anime), jujustu kaisen (anime), fullmetal alchemist (anime), demon slayer (anime), neanderthals (anthropology), pyramids of giza (archaeology), machu picchu (archaeology), stonehenge (archaeology), egyptian mummies (archaeology), giza sphinx (archaeology), notre dame cathedral (architecture), london bridge (architecture), big ben (architecture), eiffel tower (architecture), versailles (architecture), arc de triomphe (architecture), mona lisa (art), van gogh’s starry night (art), sistine chapel (art), solar eclipses (astronomy), black holes (astronomy), big bang theory (astronomy), supernovas (astronomy), dark matter (astronomy), andromeda galaxy (astronomy), elon musk vs bezos (business), taylor swift (celebrity), tom cruise (celebrity), brad pitt (celebrity), angelina jolie (celebrity), jennifer aniston (celebrity), leonardo dicaprio (celebrity), meryl streep (celebrity), robert de niro (celebrity), michael jackson (celebrity), matt damon vs jimmy kimmel (celebrity), jimmy kimmel vs jimmy fallon (celebrity), marvel vs dc (comics), batman vs superman (comics), justice league vs avengers (comics), thor vs hulk (comics), iron man vs captain america (comics), batman vs joker (comics), john constantine (comics), marvel universe (comics), dc comics (comics), big bang theory (cosmology), russian ballet (culture), bollywood (culture), hollywood (culture), alphago - the movie (documentary), the great hack (documentary), walt disney (entertainment), rose bowl parade (event), coachella (event), wimbledon (event), kentucky derby (event), rio olympics (event), tokyo olympics (event), beijing olympics (event), columbus (exploration), marco polo (exploration), nike vs adidas (fashion), world war III (fiction), pepsi vs coke (food), messi vs ronaldo (football), pele vs maradona (football), brazil vs argentina (football), monopoly (game), chess (game), go (game), dungeons and dragons (game), minecraft (game), fortnite (game), call of duty (game), mario (game), pac-man (game), sonic the hedgehog (game), fifa (game), cyberpunk 2077 (game), grand theft auto (game), the last of us (game), assassin’s creed (game), resident evil (game), halo (game), the witcher 3 (game), god of war (game), alphago vs lee sedol (games), deep blue vs kasparov (games), playstation vs xbox (gaming), mount everest (geography), grand canyon (geography), dead sea (geography), ring of fire (geography), venetian canals (geography), rocky mountains (geography), volcanic eruptions (geology), ice age (geology), mariana trench (geology), great depression (history), moon landing (history), titanic (history), viking explorers (history), stone age (history), aztec empire (history), mayan calendar (history), vikings (history), cold war (history), space race (history), moon landing (history), fall of the berlin wall (history), industrial revolution (history), edison’s light bulb (invention), the wright brothers (invention), tesla vs edison (inventors), taj mahal (landmark), great wall of china (landmark), central park (landmark), mount fuji (landmark), fifa world cup (sports), empire state building (landmark), statue of liberty (landmark), hollywood sign (landmark), golden gate bridge (landmark), niagara falls (landmark), times square (landmark), king arthur (legend), robin hood (legend), loch ness monster (legend), yeti (legend), camelot (legend), shakespearean sonnets (literature), harry potter (literature), the hobbit (literature), lord of the rings (literature), edgar allan poe (literature), charles dickens (literature), dracula (literature), frankenstein (literature), the great gatsby (literature), pride and prejudice (literature), alice in wonderland (literature), romeo and juliet (literature), moby dick (literature), war and peace (literature), little women (literature), treasure island (literature), oliver twist (literature), peter pan (literature), narnia (literature), aesop’s fables (literature), sherlock holmes (literature), the da vinci code (literature), CNN vs Fox News (media), godzilla (movie), inception (movie), matrix (movie), interstellar (movie), john wick (movie), jason bourne (movie), james bond (movie), spiderman (movie), dark knight (movie), avengers (movie), star wars (movie), indiana jones (movie), back to the future (movie), jurassic park (movie), avatar (movie), wizard of oz (movie), star wars vs star trek (movies), star wars (movies), indiana jones (movies), titanic (movies), fight club (movies), the dark knight (movies), the green mile (movies), gladiator (movies), the departed (movies), the lion king (movies), aladdin (movies), beauty and the beast (movies), little mermaid (movies), frozen (movies), tangled (movies), mulan (movies), sleeping beauty (movies), cinderella (movies), snow white (movies), the social network (movies), the beatles (music), bob dylan (music), bermuda triangle (mystery), crop circles (mystery), atlantis (myth), hercules (myth), perseus (myth), pandora’s box (myth), trojan war (myth), achilles (myth), hades (myth), olympus (myth), zeus (myth), hera (myth), apollo (myth), artemis (myth), athena (myth), poseidon (myth), phoenix (mythical creature), medusa (mythical creature), greek mythology (mythology), yellowstone (national park), yosemite (national park), grand tetons (national park), amazon rainforest (nature), sahara desert (nature), great barrier reef (nature), victoria falls (nature), niagara falls (nature), northern lights (nature), southern lights (nature), bioluminescent bays (nature), blue holes (nature), dinosaurs (paleontology), democrats vs republicans (politics), scientist vs engineer (profession), python vs c++ (programming), java vs python (programming), javascript vs java (programming), paper book vs e-book (reading), physics (science), chemistry (science), biology (science), astronomy (science), mathematics (science), albert einstein (scientist), isaac newton (scientist), charles darwin (scientist), marie curie (scientist), summer vs winter (season), twitter vs facebook (social media), olympics (sports), world cup (sports), super bowl (sports), tour de france (sports), wimbledon (sports), nba finals (sports), nfl playoffs (sports), elon musk vs mark zuckerberg (tech), ai vs human intelligence (tech), nvidia vs amd (tech), intel vs amd (tech), nvidia vs intel (tech), google vs openai (tech), iphone vs android (tech), mac vs pc (tech), ai revolution (tech), rubik’s cube (toy), barbie (toy), lego (toys), europe vs asia (travel), usa vs canada (travel), australia vs new zealand (travel), usa vs europe (travel), switzerland vs sweden (travel), beach vs mountains (vacation), city vs countryside (vacation).
Instructions: Analyze the provided song lyrics and extract the following elements: 1. Subject Matter: Write what the song is about by providing a summary of the story, narrative, central topic, or events discussed in the song lyrics. 2. Theme: Identify the main theme or message (emotion, life experience, social commentary, or philosophical concept). 3. Target Audience: Define the intended audience (age group, cultural background, or specific interests). 4. Narrative: The story and point of view (e.g., first person or second person). 5. Character Analysis: Main characters (e.g., Protagonist, Antagonist) and their traits. 6. Song Structure: Outline the structure (number of verses, choruses, bridges, intros, outros) and note any unique elements or deviations. 7. Mood: Describe the overall mood (upbeat, melancholic, introspective, etc.). 8. Reference: Identify any cultural, social, time, place, or contextual references. How to respond: You should provide your answer below after the “Answer” section. You are not allowed to use any text formatting (bold, italic, etc.) and narrative (‘Here is the answer’, ‘Below is the response’, etc.) in your answer. Only answer using the following format: subject_matter: “…..” theme: “….” target_audience: “….” narrative: “….” character_analysis: “….” song_structure: “….” mood: “….” reference: “….” Lyrics: {lyrics} Answer:
Table 7: Prompt template for extracting lyrics features (e.g subject matter, theme, mood) from real songs.
Task:
You are a talented songwriter tasked with creating a song based on the following lyrics features. The song must include all the features described below in the “Lyrics Features” section. The song should not be long; rather, it should be medium-length.
Lyrics Features:
{lyrics_feature}
Instructions:
You should write the song with metatags following the song structure from "Lyrics Features". In some very rare cases, you can also scarcely include ad libs, or non-lexical vocables.
You can add metatags to your lyrics on top of a section in [square brackets] that will create certain styles. Some examples of metatags are [Verse], [Chorus], [Bridge], [Solo], [Outro], [Pre-Chorus], [Bridge], [Hook], [Opening], [Intro], [Instrument], [Build], [Drop], [Breakdown], [Refrain], [Spoken], [Interlude], [Prelude], [Sample], etc. Adding a blank newline between sections yields the best results.
In some very rare cases, you can also scarcely use Ad libs (vocal embellishments) to your prompts in (parentheses), only when necessary. Examples include (yeah), (alright), (come on), (whoa), etc. Ad libs tend to work best at the end of a line but can also work mid-line. Unlike metatags, ad libs are sung/verbalized.
In some very rare cases, you can also scarcely use non-lexical vocables, only when necessary. Examples include la la la, na na na, sha na na.
Example:
[Verse]
I’ve been tryna call
I’ve been on my own for long enough
Maybe you can show me how to love, maybe
[Chorus]
I said, ooh, I’m blinded by the lights
No, I can’t sleep until I feel your touch
I said, ooh, I’m drowning in the night
Oh, when I’m like this, you’re the one I trust
Hey, hey, hey
Write the lyrics below after the “Lyrics:” section. You are not allowed to add any narrative or text before or after your response such as “Here’s your answer” or “Below is the response”.
Lyrics:
Table 8: Prompt template for generating song lyrics from lyrics features.
Task:
Given a song, you need to conduct a comprehensive stylistic analysis and extract all relevant information about the song’s style. This includes identifying characteristics such as instruments used (e.g., guitar, drums, piano, violin, synthesizers, bass, orchestral, solo, acoustic, trumpet, saxophone, flute, cello, harmonica, banjo, accordion, etc.), vocals types (male or female), genres (e.g., rock, pop, pop rock, indie rock, hard rock, metal, heavy metal, r&b, electronic, soul, jazz, country, reggae, classical, hip hop, blues, folk, punk, funk, disco, alternative, grunge, etc.), tempo (slow, moderate, fast), mood (e.g., melancholic, upbeat, aggressive, melodic, sad, happy, excited, nostalgic, mellow, serene, joyful, dark, gothic, etc.), and any other stylistic elements (e.g., dance, party, cinematic, dreamy, energetic, relaxing, anthemic, atmospheric, groovy, etc.) that contribute to the overall vibe, environment, or atmosphere of the song.
Your response should be structured as follows:
Answer:
<start>
style1, style2, style3, ..., styleN
<end>
For example:
Answer:
<start>
male vocals, electronic, guitar, piano, energetic, pop rock, violin, upbeat, dance, synth, sad, soul, trumpet, reggae
<end>
Please note that the list of style elements should be comprehensive and cover all relevant aspects of the song’s style. Ensure that your response follows strictly to the specified formatting, including the use of angle brackets, commas, and space separating each element written in lowercase. There must not be any narrative or text in the answer (e.g., ’Here’s your answer’ or ’Below is the response’), only the listed style elements.
{song}
Table 9: Prompt template for extracting song style (e.g. vocal type, instruments) from audio song.
Task: You are a talented songwriter, music director, and composer. Your task is to compose a {genre} genre song about {topic} with a {mood} mood. Provide the lyrics and style of the song after the “Answer” section. Follow the step-by-step instructions provided in the "Instructions" section and respond using the format given in the "How to Answer" section. Instructions: 1. Before you write the song, you need to plan what you will write about the song, then synthesize the features of the song lyrics mentioned below. Subject Matter: Write what the song is about by providing a summary of the story, narrative, central topic, or events discussed in the song lyrics. Theme: Identify the main theme or message (emotion, life experience, social commentary, or philosophical concept). Target Audience: Define the intended audience (age group, cultural background, or specific interests). Narrative: The story and point of view (e.g., first person or second person). Character Analysis: Main characters (e.g., Protagonist, Antagonist) and their traits. Song Structure: Outline the structure (number of verses, choruses, bridges, intros, and outros) and note any unique elements or deviations. Mood: Describe the overall mood (upbeat, melancholic, introspective, etc.). Reference: Identify any cultural, social, time, place, or contextual references. 2. Then, you need to write song lyrics that include all the features described in the “Lyrics Features” section. The song should not be long; rather, it should be medium-length. You should also write the lyrics with metatags following the song structure from "Lyrics Features." In some very rare cases, you can also scarcely include ad libs, or non-lexical vocables. Here are the detailed instructions: You can add metatags to your lyrics on top of a section in [square brackets] that will create certain styles. Some examples of metatags are [Verse], [Chorus], [Bridge], [Solo], [Outro], [Pre-Chorus], [Bridge], [Hook], [Opening], [Intro], [Instrument], [Build], [Drop], [Breakdown], [Refrain], [Spoken], [Interlude], [Prelude], [Sample], etc. Adding a blank newline between sections yields the best results. In some very rare cases, you can also scarcely use Ad libs (vocal embellishments) to your prompts in (parentheses), only when necessary. Examples include (yeah), (alright), (come on), (whoa), etc. Ad libs tend to work best at the end of a line but can also work mid-line. Unlike metatags, ad libs are sung/verbalized. In some very rare cases, you can also scarcely use non-lexical vocables, only when necessary. Examples include la la la, na na na, sha na na. 3. Finally, you need to compose the song by synthesizing all relevant and detailed information about the song’s style. This includes identifying characteristics such as instruments used (e.g., guitar, drums, piano, violin, synthesizers, bass, orchestral, solo, acoustic, trumpet, saxophone, flute, cello, harmonica, banjo, accordion, etc.), vocals types (female or male), genres (e.g., rock, pop, pop rock, indie rock, hard rock, metal, heavy metal, r&b, electronic, soul, jazz, country, reggae, classical, hip hop, blues, folk, punk, funk, disco, alternative, grunge, etc.), tempo (slow, moderate, fast), mood (e.g., melancholic, upbeat, aggressive, melodic, sad, happy, excited, nostalgic, mellow, serene, joyful, dark, gothic, etc.), and any other stylistic elements (e.g., dance, party, cinematic, dreamy, energetic, relaxing, anthemic, atmospheric, groovy, etc.) that contribute to the overall vibe, environment, or atmosphere of the song. The name of the style must be written in lowercase and separated by commas. 4. Finally, choose a title for the song that best suits its lyrics and style. Prompt continued on next page…
..continued from previous page How to answer: You need to provide your answer after the “Answer” section below, while strictly following the format below. Also in your answer, you are not allowed to use any text formatting (bold face, italic, etc.) or narrative (Here’s your answer, Answer is below, etc.). Just provide your answer using the below format: Lyrics Feature: <feature> * subject_matter: “…..” * theme: “….” * target_audience: “….” * narrative: “….” * character_analysis: “….” * song_structure: “….” * mood: “….” * reference: “….” </feature> Song Lyrics: <lyrics> [Verse 1] In a small town by the sea, where the waves kiss the shore, Lives a dreamer with a heart, always yearning for more. With a notebook in his hand, and a vision in his eyes, He paints the world in colors, beneath the endless skies. [Pre-Chorus] Through the struggles and the trials, he keeps his head up high, With a song within his soul, he knows he’ll touch the sky. …. </lyrics> Song Style: <style> male vocals, electronic, guitar, piano, energetic, pop rock, violin, upbeat, dance, synth, sad, soul, trumpet, reggae, …. </style> Song Title: <title> Chasing the Dream </title> Answer:
Table 10: Prompt template for generating song lyrics and style from genre, topic and mood. It also generates lyrics feature and song title as by-product.